The larger of two numbers is 8 more than four times the smaller. If the larger is increased four times the smaller, the result is 40. Find the numbers.
Answer by josmiceli(6783)

x = larger number
y = smaller number
x = 4y + 8 given
x - 4y = 8
x + 4y = 40 given, now add
2x = 48
x = 24 (the larger)
x + 4y = 40
4y = 40 - 24
4y = 16
y = 4 (the smaller)